OpenStudy (anonymous):

The functions f(x) and g(x) are described using the following equation and table: f(x) = -3(1.02)x x g(x) -1 -5 0 -3 1 -1 2 1 Which statement best compares the y-intercepts of f(x) and g(x)? The y-intercept of f(x) is equal to the y-intercept of g(x). The y-intercept of f(x) is equal to 2 times the y-intercept of g(x). The y-intercept of g(x) is equal to 2 times the y-intercept of f(x). The y-intercept of g(x) is equal to 2 plus the y-intercept of f(x).

OpenStudy (anonymous):

x= -1, 0, 1, 2 y= -5, -3, -1, 1

OpenStudy (jdoe0001):

 so to find the y-intercept for f(x), set x =0, that is \(\bf y = -3(1.02)0\) and solve for "y" to find the y-intercept for g(x), just look at the table, see when "x" is 0,

OpenStudy (anonymous):

So g(x)=-3 and f(x)=-3.06?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

so the last one is right? maybe?

OpenStudy (jdoe0001):

hmm I'd say none really, because none of the choices compare them well

OpenStudy (jdoe0001):

-3 and -3.06 do not have a "2" as multiplier, just that one of 1.06 from the parentheses

OpenStudy (anonymous):

so what one should i put for the answer

OpenStudy (jdoe0001):

hmm wait... a sec... ... darn..

OpenStudy (jdoe0001):

anyhow I guess my bad ... well... \(\bf f(x) = -3(1.02)0\implies f(x)=-3\cdot 0\implies f(x)=0\) still though

OpenStudy (jdoe0001):

unless you meant \(\bf f(x) = -3(1.02)^x\)

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

well, then recall that anything raised to 0, is 1

OpenStudy (anonymous):

Okayy

OpenStudy (jdoe0001):

\(\bf f(x) = -3(1.02)^x\implies f(x) = -3(1.02)^0\implies f(x) = -3\cdot 1 \\ \quad \\ \implies f(x) = -3\)

OpenStudy (anonymous):

So a

OpenStudy (jdoe0001):

yeap, they're both equal when x = 0 for f(x), the y-intercept is -3 from the table, for g(x) is the same

OpenStudy (anonymous):

Thank you. Can you help me with another question?

OpenStudy (jdoe0001):

you can always post anew, thus more exposure and we can revise each other

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

@hartnn could you help me with this one? i have the same table given up top but the equation i have is f(x)=-6(1.02)^x

hartnn (hartnn):

to get y-intercept of f(x) , plug in x= 0 in f(x) what do u get ?

OpenStudy (anonymous):

i got 1

hartnn (hartnn):

no... f(0) = -6(1.02)^0 = ...?

OpenStudy (anonymous):

anything to the power of zero is 1 i thought?

hartnn (hartnn):

yes, but there is -6 too! f(0) = -6*1 = -6

OpenStudy (anonymous):

oh!!!!!

OpenStudy (anonymous):

so would the answer be B?

hartnn (hartnn):

g(x) ---> -3 f(x) ----> -6 yes B :)

OpenStudy (anonymous):

can you help me in another?(=

hartnn (hartnn):

i can try :)

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

hello? @hartnn

hartnn (hartnn):

f is 5^x g is -3 f+g will be 5^x -3